OpenStudy (anonymous):

Find the nth term of each geometric sequence. a = 4, n = 7, r = -3

OpenStudy (anonymous):

Ok so we know that a sequence has the form of \[\{ a, ar, ar^2, ar^3 ...\}\] We also know that the equation when finding a specific term of a geometric sequence is \[t_n=ar^{n-1}\] So plug in the values into your equation and you will get your answer

OpenStudy (anonymous):

thank you :) so t^n=^4-^-1?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

4-1=3

OpenStudy (anonymous):

\[t_7=4(-3)^{7-1}\]

OpenStudy (anonymous):

ohh sorry okay hold on

OpenStudy (anonymous):

i typed all of that in my caculator and i got-8749

OpenStudy (anonymous):

Okkk now (-3)^6=(-3)*(-3)*(-3)*(-3)*(-3)*(-3)=729 Since there are an even amount of negative symbols then and 2 negatives make a positive so (-3)^6=729

OpenStudy (anonymous):

Then we multiply it by 4 and we get 2916

OpenStudy (anonymous):

thank you!! you are a great teacher!